The mixture contain 64g of oxygen and 60g of neon at a total pressure of 10 bar. The partial pressure in bar of oxygen and neon are

Grade:12

1 Answers

Arun
6 years ago
Dear Pratiksha
 
Mole of O2 = 64/32 = 2
Moles of Neon = 60/20 = 3
 
Mole fraction of O2 = 2/5 = 0.4
Hence mole fraction of Ne = 3/5 = 0.6
Hence partial pressure of oxygen =0.4 *10 = 4 bar
Hence partial pressure of Ne = 0.6* 10 = 6 bar
 
Regards
Arun(askIITians forum expert)